three forces of magnitude 5,4,3 units act on a particle in the direction 2i-2j-k, i+2j+2k and -2i+j-2k respectively and the particle gets displaced from point A the workdone by the forces?

Grade:12

1 Answers

Tanuja
7 years ago
Work done is product of displacement and the component of force in the direction of displacementWork done= 5(2i-2j-k)+4(I+2j+2k)+3(-2i+j-2k) = 8i+j-3k =√74